OpenStudy (anonymous):

Solve x3 = 64 over 27. ±8 over 3 8 over 3 ±4 over 3 4 over 3

OpenStudy (anonymous):

@dan815 @Nnesha

Nnesha (nnesha):

take cube root both sides to cancel out the cube

OpenStudy (anonymous):

how

Nnesha (nnesha):

here is an example \[\huge\rm \sqrt[3]{m^3} = \sqrt[3]{27}\]

OpenStudy (anonymous):

O.o

OpenStudy (anonymous):

im confuzzled

Nnesha (nnesha):

you can convert root to exponent cube root of m^3 is same as \[\huge\rm m^{3 \times \frac{ 1 }{ 3 }}\] then you can cancel out the 3's \[\huge\rm m^{\cancel{3} \times \frac{ 1 }{ \cancel{3 }}}=m\] that's how you can take cube root both sides to cancel out cube

Nnesha (nnesha):

take cube root both sides !

OpenStudy (anonymous):

*cries*

OpenStudy (anonymous):

idk how

Nnesha (nnesha):

\[\huge\rm \sqrt[3]{x^3}=\sqrt[3]{\frac{ 64 }{ 27 }}\] that's how we should take cube root cube= 3

Nnesha (nnesha):

\[\huge\rm \sqrt[3]{\frac{ a }{ b }}=\frac{ \sqrt[3]{a} }{ \sqrt[3]{b} }\]

Nnesha (nnesha):

take cube root of the numerator and take cube root of the denominator !